OpenStudy (anonymous):

Solve w + 12 = –6 – 2w.

OpenStudy (anonymous):

Bring all of the w's on one side and the constants onto the other. w + 12 = -6 - 2w 3x = -18 Solve for w now. Can you do the rest now?

OpenStudy (anonymous):

yes...ty :)

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

wat about this 1? A car salesman receives $50 for each car sale plus 2% of the purchase price. How much must a car sell for in order for the salesperson to earn at least $500?